percent: What is the yield to maturity on a 10-year, 9 percent annual coupon, $1,000 par value bond that sells for $887.002 That sells for $1,134.202 What does the fact that a bond sells at a discount or at a premium te e. (1) ll about the relationship between ra and the bonds coupon rate? you I need all parts answered using step by step. Not excel. E 1&2 and also F.
(2) What is the current yield, the capital gains yield, and the total return in each case in the preceding question? f. Suppose that the bond described in part (e) is callable in five years at a call price equal to $1,090. What is the YTC on the bond if its market value is $887? What is the YTC on the same bond if its current market price is 1 134 20
Answer #1

e)

1)

time to maturity = n = 10 years

coupon rate , c = 9% = 0.09

par value of bond , m = $1000

coupon value, x = c*m = 0.09*1000 = 90

price, p = 887

yield to maturity(ytm1) = (x+((m-p)/n))/((0.6*m)+(0.4*p)) = (90+((1000-887)/10))/((0.6*1000)+(0.4*887)) = 0.1087 or 10.87% 0r 10.9% (after rounding off)

if price , p = 1134.20

yield to maturity(ytm2) = (x+((m-p)/n))/((0.6*m)+(0.4*p)) = (90+((1000-1134.20)/10))/((0.6*1000)+(0.4*1134.20)) = 0.0708 or 7.08% (after rounding off)

when a bond sells at

a) premium , rd < coupon rate

b) discount , rd > coupon rate

2)

current yield ( for p = 887) = x/p = 90/887 = 10.15% or 10.2%

current yield ( for p = 1134.20) = x/p = 90/1134.20 = 7.94% or 7.9%

capital gains yield ( for p = 887) = ytm1 - current yield = 10.9 - 10.2 = 0.7%

capital gains yield ( for p = 1134.20) = ytm1 - current yield = 7.08 - 7.94 = -0.86%

total return ( for p = 887) = ytm1 = 10.9%

total return ( for p = 1134.20) = ytm2 = 7.08%

F)

time until bond becomes callable , n = 5

call price , m = 1090

a) price, p = 887

YTC= (x+((m-p)/n))/((0.6*m)+(0.4*p)) = (90+((1090-887)/5))/((0.6*1090)+(0.4*887)) = 0.135% or 13.5%

if price , p = 1134.20

YTC = (x+((m-p)/n))/((0.6*m)+(0.4*p)) = (90+((1090-1134.20)/5))/((0.6*1090)+(0.4*1134.20)) = 0.073 or 7.3%

Note: the above formula used to calculate YTM and YTC is an approximate formula , and hence the answers will be approximates of the real values